# Sharing is CAIRing: Characterizing Principles and Assessing Properties of Universal Privacy Evaluation for Synthetic Tabular Data

### v1

@@ Abstract
Data sharing is a necessity for innovative progress in many domains, especially in healthcare. However, the ability to share data is hindered by regulations protecting the privacy of natural persons. Synthetic tabular data provide a promising solution to address data sharing difficulties but does not inherently guarantee privacy. Still, there is a lack of agreement on appropriate methods for assessing the privacy-preserving capabilities of synthetic data, making it difficult to compare results across studies. To the best of our knowledge, this is the first work to identify properties that constitute good universal privacy evaluation metrics for synthetic tabular data. The goal of such metrics is to enable comparability across studies and to allow non-technical stakeholders to understand how privacy is protected. We identify four principles for the assessment of metrics: Comparability, Applicability, Interpretability, and Representativeness (CAIR). To quantify and rank the degree to which evaluation metrics conform to the CAIR principles, we design a rubric using a scale of 1-4. Each of the four properties is scored on four parameters, yielding 16 total dimensions. We study the applicability and usefulness of the CAIR principles and rubric by assessing a selection of metrics popular in other studies. The results provide granular insights into the strengths and weaknesses of existing metrics that not only rank the metrics but highlight areas of potential improvements. We expect that the CAIR principles will foster agreement among researchers and organizations on which universal privacy evaluation metrics are appropriate for synthetic tabular data.
@@ 1 Introduction
The volume of data collected and produced continues to increase as our society moves toward a higher degree of digitalization, and personal and health data are no exception. Regulatory restrictions on data sharing, such as HIPAA [15] and GDPR [7], have been implemented to protect individuals from privacy violations. The downside of imposing restrictions on data sharing is the substantial limitations on innovation and research based on those data. Consequently, methods that allow data sharing under the regulations are a necessity to catalyze data-sharing capabilities and privacy improvements. Privacy-preserving synthetic data generation is a promising attempt to solve this problem by generating simulated records with the same statistical properties as the true data [18, 17]. However, as synthetic data generation is still maturing for tabular data, an abundance of evaluation metrics are applied in various studies. Consequently, comparing the evaluation results of the generated data is a highly challenging task, which slows innovation in generation methods [2, 8, 11]. Therefore, standardized approaches are needed to evaluate the quality of synthetic data. In addition, more regulations are in the works as of writing this paper, namely the European Union Artificial Intelligence Act [6] which moves towards a more secure, fair, and private use of artificial intelligence (AI). Therefore, privacy evaluation is important when dealing with synthetic data. Many studies claim that fully synthetic data are inherently private [11, 14] due to the disconnect between true and synthetic data. However, it has been shown that synthetic data can be successfully attacked and therefore require a thorough evaluation of privacy [12, 4].
When we combine the knowledge that 1) synthetic data are promising, 2) there is no consensus for evaluation metrics, 3) regulations are focusing more on privacy in AI, and 4) synthetic data are not inherently private; it is clear that we need to define how we best assess the privacy levels provided by synthetic data. The goal is to allow for comparability across works in a way that allows non-technical stakeholders, such as regulators and natural persons, to understand how privacy is protected.
@@
Therefore, the aim of this paper is to define principles for good universal privacy evaluation metrics for synthetically generated tabular data that can aid researchers and regulators in creating and selecting appropriate evaluation methods. We have created a scoring system that uses a rubric to rank existing evaluation metrics in order to achieve that goal. In this work, we use the concept of privacy metric and privacy measure synonymously for any function, method, algorithm, or framework that is used to evaluate the privacy level provided by synthetically generating data. To the best of our knowledge, this is the first attempt to define good practices with a scoring system for the evaluation of the privacy of synthetic tabular data.
@@
In the remainder, we present related work (Section 2), followed by a presentation of the criteria and rubric developed in this work (Section 3). Then we detail all dimensions of the rubric (Section 4) and apply them to a selection of existing privacy evaluation metrics (Section 5). Lastly we discuss the strengths, limitations, and significance of our work (Section 6).
@@ 2 Related Work
One of the primary motivations for creating synthetic data is to introduce privacy, but the evaluation of utility and realism has been the main focus of many researchers [14]. The same focus on utility is also present when defining universally applicable evaluation metrics, with most research revolving around finding suitable utility or realism metrics. El Emam et al. [5] attempt to quantify how well six different utility metrics can rank the performance of synthetic data generation methods according to a logistic regression prediction. They empirically evaluate 30 different datasets using three SDG methods. The baseline against which they test is the difference between prediction performance for the true and synthetic data using the area under the receiver operating characteristic curve and the area under the precision-recall curve. Using the Page test, they test whether the utility metrics agree with the baseline. This approach enables empirical quantification of the metrics’ performances but is limited to the single scenario of using logistic regression. Dankar et al. [3] identify four classes of utility metrics for which they select a representative utility metric that they test. Ultimately, they assess whether one of the metrics can be used for overall utility assessment by investigating the correlations between the selected metrics. The authors identify the Hellinger distance as somewhat correlated with two of the three metrics and conclude that it can be used as a standardized utility metric. The study focuses on a limited set of criteria for streamlining evaluation of utility and does not consider elements such as attribute heterogeneity. Hernandez et al. [10, 11] identify the issue of comparing evaluation results across studies. Therefore, they opt to use a proxy scale consisting of “excellent”, “good”, and “poor” and define the criteria to fall into each level for all included evaluation metrics. The downside of using this scale is the limited granularity for interpretation. Furthermore, not all metrics fit naturally into the three levels. For example, they distribute the result for membership inference attacks (MIAs) [18] into “excellent” if privacy is preserved and “good” otherwise. Consequently, even though that method is a step in the right direction for comparability, it is non-exhaustive and provides a limited proxy for evaluation.
@@ 3 Method
Assessing the quality of complex work such as privacy evaluation metrics can be a challenging task to perform reproducibly and objectively. Therefore, we break down the assessment into smaller dimensions using a rubric. Rubrics have been successful in educational settings in terms of assessment of lecturer-student, student-lecturer, and peer assessments [1], but also provide an important assessment tool in other fields of assessment, including privacy [16]. Each dimension should be defined on a scale where the requirements to obtain a specific score are clearly defined to avoid subjectivity and bias. For this work, four general categories have been identified based on the ideal of a universal privacy evaluation metric and the stakeholders of it:
@@
Comparability: The ideal metric is comparable across different synthetic datasets.
@@
Applicability: The ideal metric is applicable to all datasets, data types, and SDG methods.
@@
Interpretability: The ideal metric can be easily communicated and understood unambiguously by stakeholders.
@@
Representativeness: The ideal metric accurately represents a realistic privacy level.
Figure 1: The four CAIR principles with their respective dimensions.
@@
The four categories form the acronym CAIR and represent the fact that a universal metric for tabular synthetic data should be usable for all data (C and A) and trustworthy (I and R). Each category is further broken down into four dimensions, making the rubric consist of 16 total assessments (illustrated in Figure 1). Having the same number of unweighted dimensions for each category ensures equal weighting for the general categories as well. The final result is an unweighted mean of all 16 scores with the standard error. For the sake of accuracy, we allow the inclusion of .5 decimals if a metric is deemed to fall between two scores. The complete rubric is presented in Table 1 using a scale of 1-4 (poor-excellent) with a higher score preferred. Details of each dimension are presented in Section 4.

### v2

@@ Abstract
Data sharing is a necessity for innovative progress in many domains, especially in healthcare. However, the ability to share data is hindered by regulations protecting the privacy of natural persons. Synthetic tabular data provide a promising solution to address data sharing difficulties but does not inherently guarantee privacy. Still, there is a lack of agreement on appropriate methods for assessing the privacy-preserving capabilities of synthetic data, making it difficult to compare results across studies. To the best of our knowledge, this is the first work to identify properties that constitute good universal privacy evaluation metrics for synthetic tabular data. The goal of universally applicable metrics is to enable comparability across studies and to allow non-technical stakeholders to understand how privacy is protected. We identify four principles for the assessment of metrics: Comparability, Applicability, Interpretability, and Representativeness (CAIR). To quantify and rank the degree to which evaluation metrics conform to the CAIR principles, we design a rubric using a scale of 1-4. Each of the four properties is scored on four parameters, yielding 16 total dimensions. We study the applicability and usefulness of the CAIR principles and rubric by assessing a selection of metrics popular in other studies. The results provide granular insights into the strengths and weaknesses of existing metrics that not only rank the metrics but highlight areas of potential improvements. We expect that the CAIR principles will foster agreement among researchers and organizations on which universal privacy evaluation metrics are appropriate for synthetic tabular data.
keywords:
Privacy , Privacy evaluation , Synthetic data , Data sharing , Tabular data , Machine learning
journal: Machine Learning with Applications
[…]
[1]organization=University of Southern Denmark, Department of Mathematics and Computer Science, addressline=Campusvej 55, city=Odense, postcode=5230, country=Denmark
22footnotetext: Corresponding author.
@@ 1 Introduction
The volume of data collected and produced continues to increase as our society moves toward a higher degree of digitalization, and personal and health data are no exception. Regulatory restrictions on data sharing, such as HIPAA (Office for Civil Rights , 2012, OCR) and GDPR (European Parliament and Council of the European Union,, 2016), have been implemented to protect individuals from privacy violations. The downside of imposing restrictions on data sharing is the substantial limitations on innovation and research based on these data. Consequently, methods that allow data sharing under the regulations are a necessity to catalyze data-sharing capabilities and privacy improvements. Privacy-preserving synthetic data generation (SDG) is a promising attempt to solve this problem by generating simulated records with the same statistical properties as the true data (Rankin et al.,, 2020; Sun et al.,, 2021). Synthetic records can be used as a proxy for real data, while allowing analyses to reach highly similar conclusions, thereby maintaining the privacy of individuals. However, no standardized approach to evaluating the privacy preserving capabilities of synthetic data has been established. Consequently, studies employ a variety of incomparable measures that make cross-study comparisons challenging. This work aims at establishing a foundation on which universally applicable metrics can be developed and improved to consolidate the research community and regulatory units.
1.1 Privacy Preserving Methods
Many privacy preserving technologies have been proposed to ensure the protection of individuals. Traditional anonymization techoniques such as k-Anonymity (Samarati and Sweeney,, 1998; Sweeney,, 2002) and $\ell$-diversity (Machanavajjhala et al.,, 2006) protect privacy by generalizing records to ensure that no unique individuals are present in the data. These methods are simple but have been shown to be vulnerable to adversarial attacks (Narayanan and Shmatikov,, 2008; Li et al.,, 2007). A different approach is Fully Homomorphic Encryption (FHE) which allows analyses to be conducted on a cipher text rather the original data (Gentry,, 2009). However, FHE is computationally expensive, making it impractical for many use cases such as deep learning (Lee et al.,, 2022). A popular method to introduce privacy in algorithms is Differential Privacy (DP), which provides formal privacy guarantees to an algorithm by adding noise to the data or employing noisy operations (Dwork and Roth,, 2013). Federated learning is a privacy-oriented method to address machine learning problems in which sensitive data are located at multiple clients (Zhang et al.,, 2021). In a federated learning setting, machine learning models are trained locally and only the model parameters are transferred to a central unit that aggregates models from all clients to a single global model.
These methods all have their specific use cases and can often be used in combination with each other such as training differentially private models in a federated learning setting. Similarly, synthetic data is useful for specific use-cases and be created using other privacy preserving methods. In cases where the downstream tasks are unknown or where data sharing is needed, synthetic data is a promising solution as an alternative to other privacy preserving methods (Rankin et al.,, 2020; Sun et al.,, 2021).
1.2 Importance of Privacy Evaluation
As synthetic data generation is still maturing for tabular data, an abundance of evaluation metrics is applied in various studies. Consequently, comparing the evaluation results of the generated data is a highly challenging task, slowing innovation in generation methods (Chundawat et al.,, 2022; Ghosheh et al.,, 2022; Hernandez et al.,, 2022). Therefore, standardized approaches are needed to evaluate the privacy-preserving capabilities of synthetic tabular data.
In addition, the European Union address potential issues with the application of artificial intelligence (AI) with the Artificial Intelligence Act (European Comission,, 2021) which moves towards a more secure, fair, and private use of AI. Therefore, privacy evaluation is important when dealing with synthetic data. Multiple studies claim that fully synthetic data are inherently private (Hernandez et al.,, 2022; Murtaza et al.,, 2023) due to the disconnect between true and synthetic data. However, synthetic data have been shown to be successfully attacked and therefore require a thorough privacy evaluation (El Emam et al.,, 2020; Hittmeir et al.,, 2020).
1.3 Privacy Evaluation
Many methods have been proposed to assess the privacy preserving capabilities of synthetic tabular data (van Breugel et al.,, 2023; D’Amico et al.,, 2023; El Emam et al.,, 2020; Hu et al.,, 2023; Rashidian et al.,, 2020; Tucker et al.,, 2020; Yale et al.,, 2019; Yoon et al.,, 2020) all of which serve as important steps in ensuring privacy preservation. However, parallel to the development of privacy evaluation metrics is the development of generation methods to establish privacy (Hernandez et al.,, 2022; Murtaza et al.,, 2023). Nevertheless, when a wide selection of evaluation metrics is available at evaluation time, studies that develop generation methods can select different metrics, causing difficulties in comparing results between works. This effect is exacerbated by evaluation metrics being designed for specific purposes or data types making them inapplicable in some cases.
Although the purpose of privacy evaluation metrics is to accurately gauge the privacy level provided by the generation process, research may be slowed down by having incomparable results across studies. Therefore, there is a need for universal privacy evaluation metrics for synthetic tabular data that can be applied to and compared between multiple studies. Considering the important role regulations have in protecting privacy, we argue that a good universal evaluation metric allows for comparability across works in a way that allows non-technical stakeholders, such as regulators and natural persons, to understand how privacy is protected.
1.4 Main Contributions
@@
Therefore, the objective of this paper is to define principles for good universal privacy evaluation metrics for synthetically generated tabular data that can aid researchers and regulators in creating and selecting appropriate evaluation methods. We have created a scoring system that uses a rubric to rank existing evaluation metrics on their universality to achieve that goal. A byproduct of establishing universally applicable metrics is that the scoring is not designed to rank metrics by their accuracy, but rather by their versatility across different applications while still maintaining precision in estimating privacy levels.
The main contributions of our work are the following:
  • 1.
    Characterization of four general principles for universal privacy evaluation methods for tabular synthetic data named CAIR (Comparability, Applicability, Interpretability, and Representativeness). Each principle consists of four dimensions that provide high-fidelity insight into the universality of privacy evaluation metrics.
  • 2.
    We provide a scoring rubric for CAIR based on the identified principles that enable users to quantify the degree of conformity to CAIR. The scoring allows granular insights while mitigating bias from qualitatively scoring metrics.
In this work, we use the concept of privacy metric and privacy measure synonymously for any function, method, algorithm, or framework that is used to evaluate the privacy level provided by synthetically generating data. To the best of our knowledge, this is the first attempt to define good practices with a scoring system for the evaluation of the privacy of synthetic tabular data.
@@
In the remainder, we first present related work, followed by a presentation of the criteria and rubric developed in this work. Then we detail all dimensions of the rubric and apply them to a selection of existing privacy evaluation metrics. Finally, we discuss the strengths, limitations, and significance of our work followed by a conclusion. As a courtesy for the reader, we supply additional details in A.
@@ 2 Related Work
One of the primary motivations for creating synthetic data is to introduce privacy, but the evaluation of utility and realism has been a primary focus of many researchers (Baowaly et al.,, 2019; van Breugel et al.,, 2023; Chandra et al.,, 2022; Chundawat et al.,, 2022; Dankar et al.,, 2022; Farou et al.,, 2020; Ghosheh et al.,, 2022; Hansen et al.,, 2023; Murtaza et al.,, 2023; Rankin et al.,, 2020). The same focus on utility is also present when defining universally applicable evaluation metrics, with most research focusing on finding suitable utility or realism metrics (Chundawat et al.,, 2022; Dankar et al.,, 2022; El Emam et al.,, 2022; Hernandez et al.,, 2022; Hernadez et al.,, 2023).
2.1 Privacy Evaluation Approaches
Among the many approaches to privacy evaluation are three broad categories (Ganev and Cristofaro,, 2023; Murtaza et al.,, 2023): attribute disclosure, membership inference, and similarity-based evaluation.
Attribute disclosure risk (ADR) is an attack that evaluates an adversary’s capability of predicting previously unknown sensitive attributes for individuals, with various levels of adversarial knowledge (Kaur et al.,, 2021).
Membership inference attacks (MIA) aim to infer whether an individual is in the training data. Many variations of MIAs exist ranging from simple attacks based on the proximity of synthetic records to a real record (Yan et al.,, 2020) to more advanced attacks that leverage underlying structures in synthetic and real data such as DOMIAS (van Breugel et al.,, 2023), LOGAN (Hayes et al.,, 2017), ReconSyn (Ganev and Cristofaro,, 2023), and privacy gain defined by Stadler et al., (2022).
Similarity-based metrics use similarity or proximity between synthetic and real records to determine privacy violations. There is a wide range of similarity-based metrics, from the relatively straightforward Distance to Closest Record (DCR) (Guillaudeux et al.,, 2023) to more complex metrics such as $\epsilon$-identifiability (Yoon et al.,, 2020) and identity disclosure risk (IDR) by El Emam et al., (2020).
Across all these methods, covering the three categories, there is a shared goal: optimizing the accuracy of privacy evaluation. Although striving for accuracy is crucial, the practical adoption of these metrics faces challenges when they are incompatible with diverse scenarios.
2.2 Addressing Comparability Challenges
El Emam et al., (2022) attempt to quantify how well six different utility metrics can rank the performance of synthetic data generation methods according to a logistic regression prediction. They empirically evaluate 30 different datasets using three SDG methods. The baseline against which they test is the difference between prediction performance for the true and synthetic data using the area under the receiver operating characteristic curve and the area under the precision-recall curve. Using the Page test (Siegel and Castellan,, 1988), they test whether the utility metrics agree with the baseline. This approach enables empirical quantification of the metrics’ performances but is limited to the single scenario of using logistic regression.
Dankar et al., (2022) identify four classes of utility metrics for which they select a representative utility metric for testing. Ultimately, they assess whether one of the metrics can be used for overall utility assessment by investigating the correlations between the selected metrics. The authors identify the Hellinger distance as somewhat correlated with two of the three metrics and conclude that it can be used as a standardized utility metric. The study focuses on a limited set of criteria for streamlining evaluation of utility and does not consider elements such as attribute heterogeneity.
Hernandez et al., (2022); Hernadez et al., (2023) identify challenges comparing results across studies. Therefore, they use a proxy scale consisting of “excellent”, “good”, and “poor” and define the criteria to fall into each level for all included evaluation metrics. The downside of using this scale is the limited granularity for interpretation. Furthermore, not all metrics fit naturally into the three levels. For example, they distribute the result for membership inference attacks (MIAs) (Sun et al.,, 2021) into “excellent” if privacy is preserved and “good” otherwise. Thus, even though the method is a step in the right direction for comparability, it provides a limited proxy for utility and privacy evaluation.
The works presented above design methods that emphasize the utility over privacy to compare results across studies. Furthermore, the approaches have limitations in terms of granularity and applicability, causing the methods to function in a narrow field. Certainly, any attempt at addressing comparability challenges is a step in a constructive direction, but further standardization is required to allow for universally applicable evaluation metrics. Therefore, this work aims to provide a foundation on which universal privacy evaluation metrics can be developed and improved with a common direction. Ideally, such a foundation can consolidate researchers and regulatory units and form agreement for future regulations. Furthermore, addressing these comparability issues can facilitate innovative acceleration by allowing researchers to compare results between works, improving informed decision making when designing new privacy-aware generative approaches.
Figure 1: The four CAIR principles with their respective dimensions.
@@ 3 Method
Assessing the quality of complex work such as privacy evaluation metrics can be a challenging task to perform reproducibly and objectively. Therefore, we break down the assessment into smaller dimensions using a rubric. Rubrics have been successful in educational settings in terms of assessment of lecturer-student, student-lecturer, and peer assessments (Bowen-Mendoza et al.,, 2022), but also provide an important assessment tool in other fields of assessment (Lautrup et al.,, 2023; O’Donnell et al.,, 2022). Each dimension should be defined on a scale where the requirements to obtain a specific score are clearly defined to avoid subjectivity and bias.
For this work, four general principles have been identified based on the ideal of a universal privacy evaluation metric and the stakeholders of it:
@@
Comparability: The ideal privacy metric is comparable across different synthetic datasets.
@@
Applicability: The ideal privacy metric is applicable to all datasets, data types, and SDG methods.
@@
Interpretability: The ideal privacy metric can be easily communicated and understood unambiguously by stakeholders.
@@
Representativeness: The ideal privacy metric accurately represents a realistic privacy level.
@@
The four principles form the acronym CAIR and represent the fact that a universal metric for tabular synthetic data should be usable for all data (Comparability and Applicability) and trustworthy (Interpretability and Representativeness). Each principle is further broken down into four dimensions, which makes the rubric consist of 16 total assessments as illustrated in Figure 1. Maintaining an equal number of unweighted dimensions across all categories ensures uniform weighting among the general categories. The final result is an unweighted mean of all 16 scores. For the sake of accuracy, we allow the inclusion of .5 decimals if a metric is deemed to fall between two scores. The complete rubric is presented in Table LABEL:tab:rubric using a scale of 1-4 (poor-excellent) with a higher score preferred. Details of each dimension are presented in Section 4.
Metrics can be scored either by a single evaluator or through a multi-evaluator approach. Both settings carry benefits, as the single evaluator setting allows for group discussions and quick estimates of improvement potentials of a metric while the multi-evaluator setting aims at minimizing qualitative biases.
3.1 Single Evaluator
In a single evaluator setting, an individual scores a metric on the 16 CAIR dimensions. Using this setting is sensitive to potential evaluator bias but serves an important purpose when designing or improving new privacy evaluation metrics. By allowing for single evaluators, researchers can obtain granular insights into strengths and weaknesses of a metric in development without the need to consult independent evaluators.
The overall CAIR score is given by the mean of all the dimensions. Let $N$ be the number of CAIR dimensions (16 in this work). Let $d_{i}$ denote the $i^{th}$ dimensions such that $d_{i}\in[1,1.5,\dots,4]$. Then, the CAIR score for a single evaluator is given by:
$\operatorname{CAIR}_{single}=\frac{1}{N}\sum_{i=1}^{N}d_{i}.$ (1)
To encourage discussions during metric development, a group of people can collectively assume the role of a single evaluator and report a single score for each dimension.
3.2 Multiple Evaluators
To mitigate qualitative bias by individual evaluators, multiple independent evaluators can score each metric on the 16 dimensions. Contrary to the single-evaluator setting of a group, this setting requires all evaluators to individually score the given metric. The standard error is used as the uncertainty estimate based on the assumption that by increasing the number of evaluators, the CAIR score approaches a consensus. Consequently, the error must be propagated to the final CAIR score. The evaluation workflow is depicted in Figure 2 and is defined below.
To account for multiple evaluators, we extend Equation 1 accordingly. Let $N$ be the number of CAIR dimensions. Let $\mathcal{E}$ be a domain of individual independent evaluators $E\in\mathcal{E}$, i.e., functions such that $E:d_{i}\rightarrow\{1,1.5,\dots,4\}$, $i=1\dots N$ where $d_{i}$ is the corresponding CAIR dimension. Then, the CAIR score is defined as:
$\operatorname{CAIR}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathcal{E}|}\sum_{E\in\mathcal{E}}E(d_{i}).$ (2)
[…]
Figure 2: Visual representation of the multi-evaluator CAIR evaluation workflow defined in Equation 2. Given the definition of a privacy metric, a series of independent evaluators score the metric on the CAIR dimensions. For each dimension, the mean across the evaluators’ scores are computed with the final CAIR score being the average of dimension means.
3.3 Levels of Granularity
Segmenting the principles into multiple dimensions allows for varying levels of granularity when examining the assessments. Figure 3 presents three levels of granularity, each with different benefits. The overall CAIR score in Figure 3(a) is the most general representation, being an aggregate of all dimensions across all evaluators. It provides a single value, that is easy to compare between metrics. Figure 3(b) groups the scores according to the respective principles and provides internal propagated errors. This representation allows differentiation in performance across principles rather than over all dimensions. The most detailed representation is presented in Figure 3(c) where the mean score for each dimension across evaluators. This representation enables a detailed investigation into strengths and weaknesses. For the most comprehensive investigation, the dimension-wise representation in Figure 3(c) is recommended.
(a) CAIR Score
(b) Grouped by Principles
(c) Individual Dimensions
Figure 3: An illustrative example of CAIR assessments displayed with varying levels of granularity. (a) displays the general CAIR score as an aggregate value. (b) presents the scores grouped by the corresponding principle. (c) illustrates scores on individual dimensions.
4 The CAIR Principles
The CAIR principles for good privacy evaluation for synthetic tabular data consist of a range of criteria that aim to describe the ideal privacy measure. The rubric in Table LABEL:tab:rubric is limited in description depth, but should be sufficient to score any privacy metric without additional information. However, more detailed information on the dimensions is necessary to adequately argue for their inclusion.
4.1 Comparability
One of the primary motivations for advocating standardized privacy evaluation metrics lies in the concept of comparability. By establishing a consistent set of metrics, researchers and practitioners can systematically compare privacy outcomes across various models, datasets, and projects. This comparability not only facilitates informed decision-making but also fosters healthy competition within the research community of synthetic data generation. With the ability to assess the efficacy of different generation methods, stakeholders can identify the most effective strategies for preserving privacy. Comparisons accelerate innovation by allowing results and generation methods to be compared and shared between domains and context-specific applications.
C1 Scale
Comparisons of privacy levels between generation processes rely on the privacy measure being on the same scale. If the scale is not identical, the resulting evaluation values are incompatible, making any attempt to compare the results futile. An argument can be made that users should normalize the data before computing the evaluation results. However, considering that synthetically generated data are primarily useful on the same scale as the true data, it is unlikely that all users will report the results on normalized data. In addition, the scale of the data can skew the results of a measure even if the metric has well-defined bounds. Consider the function $f(\mathcal{T},\mathcal{S})=\sigma(\lVert\mathcal{T}_{i}-\mathcal{S}_{j}\rVert_{2})$, where $\mathcal{T}$ and $\mathcal{S}$ are true and synthetic data respectively and $\sigma$ is the sigmoid function; then if two attributes defined as $a_{0}\in(0,10^{6})$ and $a_{1}\in(0,1)$ are in the dataset, the influence of $a_{0}$ on the function will be substantially greater than that of $a_{1}$ even though the final result is in $(0,1)$.
C2 Metric Bounds
C3 Data Type Agnostic
Heterogeneous data types are prevalent in tabular data (Murtaza et al.,, 2023). As such, it is crucial that the heterogeneity of the data has a limited effect on the output of the metric. Ideally, the metric should be invariant to any tabular data type, but a remediation can be to perform appropriate pre-processing. If pre-processing cannot mitigate adverse effects of the mixed data types, further analysis should be conducted to identify whether a systematic bias exists towards any specific subsets of data types. Consider the use of Hamming distance to determine the similarity between synthetic and true data. Hamming distance computes the number of attributes where the records differ and as such performs best for categorical data. Some pre-processing such as data binning can mitigate parts of the inaccuracies on continuous data, but it can potentially have significant adverse effects on the results.
C4 Cross-Domain Relevance